# Equal contribution    Project lead    Corresponding author

### v3

@@ 1 Introduction
Time series analysis is crucial for identifying and predicting temporal patterns across various domains, including weather forecasting (Bi et al., 2023), medical symptom classification (Kiyasseh et al., 2021), anomaly detection in spacecraft monitoring (Xu, 2021), and imputing missing data in wearable sensors (Wu et al., 2020). These diverse applications highlight the versatility and importance of time series analysis in addressing real-world challenges. A key advancement in this field is the development of time series pattern machines (TSPMs), which aim to create a unified model architecture capable of handling a broad range of time series tasks across domains (Zhou et al., 2023; Wu et al., 2023).
@@ 4.1.7 Classification and Anomaly detection
Figure 3: Results of classification and anomaly detection. The results are averaged from several datasets. Higher accuracy and F1 score indicate better performance. $\ast$. indicates the Transformer-based models. See Table 21 and 22 in the Appendix H for full results.
@@ Appendix A Implementation Details
To assess the effectiveness of our method across various tasks, we select 27 advanced baseline models spanning a wide range of architectures. Specifically, we utilize CNN-based models: MICN (2023a), SCINet (2022a), and TimesNet (2023); MLP-based models: TimeMixer (2024b), LightTS (2022a), and DLinear (2023); RMLP&RLinear (2023a) and Transformer-based models: iTransformer (2024), PatchTST (2023), Crossformer (2023), FEDformer (2022b), Stationary (2022b), Autoformer (2021), and Informer (2021b). These models have demonstrated superior capabilities in temporal modeling and provide a robust framework for comparative analysis. For specific tasks, TiDE (2023a), FiLM (2022a), N-HiTS (2023), and N-BEATS (2019) address long- or short-term forecasting; Anomaly Transformer (2021) and MTS-Mixers (2023b) target anomaly detection; while Rocket (2023a), LSTNet (2018c), LSSL (2021), and Flowformer (2022) are utilized for classification. Few/zero-shot forecasting tasks employ ETSformer (2022), Reformer (2020), and LLMTime (2023).
@@ Appendix E Efficiency Analysis
We comprehensively compare the forecasting and imputation in performance, training speed, and memory footprint of the following models: TimeMixer++, iTransformer(Liu et al., 2024), PatchTST(Nie et al., 2023), TimeMixer(Wang et al., 2024b), TIDE(Das et al., 2023a), Fedformer(Zhou et al., 2022b), TimesNet(Wu et al., 2023), MICN(Wang et al., 2023a), and SCINet(Liu et al., 2022a).
@@ Appendix G Error bars
Solar-Energy $0.203\pm 0.001$ $0.238\pm 0.010$ $0.233\pm 0.009$ $0.262\pm 0.007$ 99%
Electricity $0.165\pm 0.011$ $0.253\pm 0.001$ $0.178\pm 0.002$ $0.270\pm 0.005$ 99%
Traffic $0.416\pm 0.015$ $0.264\pm 0.013$ $0.428\pm 0.008$ $0.282\pm 0.002$ 99%
ETTh1 $0.419\pm 0.011$ $0.432\pm 0.015$ $0.454\pm 0.004$ $0.447\pm 0.007$ 99%
ETTh2 $0.339\pm 0.009$ $0.380\pm 0.002$ $0.383\pm 0.004$ $0.407\pm 0.007$ 99%
ETTm1 $0.369\pm 0.005$ $0.378\pm 0.007$ $0.407\pm 0.004$ $0.410\pm 0.009$ 99%
ETTm2 $0.269\pm 0.002$ $0.320\pm 0.012$ $0.288\pm 0.010$ $0.332\pm 0.003$ 99%
@@
Dataset MAE MAPE RMSE MAE MAPE RMSE Interval
@@
Dataset SMAPE MASE OWA SMAPE MASE OWA Interval
@@ Appendix H Full Results
Due to the space limitation of the main text, we place the full results of all experiments in the following: long-term forecasting in Table 16, univariate short-term forecasting in Table 17, multivariate short-term forecasting in Table 18, imputation in Table 19, few-shot learning in Table 20, classification in Table 21 and anomaly detection in Table 22.
@@
Table 19: Full results of imputation task across six datasets. To evaluate our model performance, we randomly mask $\{12.5\%,25\%,37.5\%,50\%\}$ of the time points in time series of length 1024. The final results are averaged across these 4 different masking ratios.
Models
TimeMixer++
TimeMixer
iTransformer
PatchTST
Crossformer
FEDformer
TIDE
DLinear
TimesNet
MICN
Autoformer
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
ETTm1
0.041
0.127
0.072
0.178
0.075
0.177
0.097
0.194
0.0810
0.196
0.048
0.152
0.090
0.210
0.076
0.191
0.049
0.147
0.059
0.170
0.062
0.169
ETTm2
0.024
0.135
0.061
0.166
0.055
0.169
0.080
0.183
0.111
0.219
0.087
0.198
0.169
0.263
0.088
0.198
0.035
0.124
0.109
0.221
0.077
0.181
ETTh1
0.091
0.198
0.152
0.242
0.130
0.213
0.178
0.231
0.201
0.309
0.099 0.225
0.289
0.395
0.174
0.288
0.142
0.258
0.127
0.254
0.148
0.267
ETTh2
0.065
0.157
0.101
0.294
0.125
0.259
0.124
0.293
0.206
0.308
0.262
0.344
0.709
0.596
0.120
0.238
0.088 0.198
0.179
0.290
0.129
0.242
ECL
0.109
0.197
0.142
0.261
0.140
0.223
0.129
0.198 0.125
0.204
0.181
0.314
0.182
0.202
0.080
0.200
0.135
0.255
0.138
0.246
0.141
0.234
Weather
0.049
0.078
0.091
0.114
0.095
0.102
0.082
0.149
0.150
0.111
0.064
0.139
0.063
0.131
0.071
0.107
0.061 0.098
0.075
0.126
0.066
0.107
Models
TimeMixer++
TimeMixer
iTransformer
TiDE
Crossformer
DLinear
PatchTST
TimesNet
FEDformer
Autoformer
Stationary
ETSformer
LightTS
Informer
Reformer
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
ETTh1
0.517
0.512
0.613
0.52
0.51
0.597
0.589
0.535
0.645
0.558
0.691
0.600
0.633
0.542
0.869
0.628
0.639
0.561
0.702
0.596
0.915
0.639
1.180
0.834
1.375
0.877
1.199
0.809
1.249
0.833
ETTh2
0.379
0.391
0.402
0.433
0.455
0.461
0.395 0.412
0.428
0.447
0.605
0.538
0.415
0.431
0.479
0.465
0.466
0.475
0.488
0.499
0.462
0.455
0.894
0.713
2.655
1.160
3.872
1.513
3.485
1.486
ETTm1
0.398
0.431
0.487
0.461
0.491
0.516
0.425
0.458
0.462
0.489
0.411
0.429
0.501
0.466
0.677
0.537
0.722
0.605
0.802
0.628
0.797
0.578
0.980
0.714
0.971
0.705
1.192
0.821
1.426
0.856
ETTm2
0.291
0.351
0.311
0.367
0.375
0.412
0.317
0.371
0.343
0.389
0.316
0.368
0.296
0.343
0.320
0.353
0.463
0.488
1.342
0.930
0.332
0.366
0.447
0.487
0.987
0.756
3.370
1.440
3.978
1.587
Weather
0.241
0.271
0.242
0.281
0.291
0.331
0.249
0.291
0.267
0.306
0.241
0.283
0.242
0.279
0.279
0.301
0.284
0.324
0.300
0.342
0.318
0.323
0.318
0.360
0.289
0.322
0.597
0.495
0.546
0.469
ECL
0.168
0.271
0.187
0.277
0.241
0.337
0.196
0.289
0.214
0.308
0.180
0.280
0.180 0.273
0.323
0.392
0.346
0.427
0.431
0.478
0.444
0.480
0.660
0.617
0.441
0.489
1.195
0.891
0.965
0.768
@@
  • $\ast$ The original paper of Anomaly Transformer (Xu, 2021) adopts the temporal association and reconstruction error as a joint anomaly criterion. For fair comparisons, we only use reconstruction error here.

### v5

@@ 1 Introduction
Time series analysis is crucial for identifying and predicting temporal patterns across various domains, including weather forecasting (Bi et al., 2023), medical symptom classification (Kiyasseh et al., 2021), anomaly detection in spacecraft monitoring (Xu et al., 2022), and imputing missing data in wearable sensors (Wu et al., 2020). These diverse applications highlight the versatility and importance of time series analysis in addressing real-world challenges. A key advancement in this field is the development of time series pattern machines (TSPMs), which aim to create a unified model architecture capable of handling a broad range of time series tasks across domains (Zhou et al., 2023; Wu et al., 2023).
@@ 4.1.7 Classification and Anomaly detection
Figure 3: Results of classification and anomaly detection. The results are averaged from several datasets. Higher accuracy and F1 score indicate better performance. $\ast$. indicates the Transformer-based models. See Table 19 and 20 in the Appendix H for full results.
@@ Appendix A Implementation Details
To assess the effectiveness of our method across various tasks, we select 27 advanced baseline models spanning a wide range of architectures. Specifically, we utilize CNN-based models: MICN (2023a), SCINet (2022a), and TimesNet (2023); MLP-based models: TimeMixer (2024b), LightTS (2022a), and DLinear (2023); RMLP&RLinear (2023a) and Transformer-based models: iTransformer (2024), PatchTST (2023), Crossformer (2023), FEDformer (2022b), Stationary (2022c), Autoformer (2021), and Informer (2021b). These models have demonstrated superior capabilities in temporal modeling and provide a robust framework for comparative analysis. For specific tasks, TiDE (2023b), FiLM (2022a), N-HiTS (2023), and N-BEATS (2019) address long- or short-term forecasting; Anomaly Transformer (2022) and MTS-Mixers (2023b) target anomaly detection; while Rocket (2023a), LSTNet (2018c), LSSL (2022a), and Flowformer (2022) are utilized for classification. Few/zero-shot forecasting tasks employ ETSformer (2022), Reformer (2020), and LLMTime (2023).
@@ Appendix E Efficiency Analysis
We comprehensively compare the forecasting and imputation in performance, training speed, and memory footprint of the following models: TimeMixer++, iTransformer(Liu et al., 2024), PatchTST(Nie et al., 2023), TimeMixer(Wang et al., 2024b), TIDE(Das et al., 2023b), Fedformer(Zhou et al., 2022b), TimesNet(Wu et al., 2023), MICN(Wang et al., 2023a), and SCINet(Liu et al., 2022a).
@@ Appendix G Error bars
Solar-Energy $0.203\pm 0.027$ $0.238\pm 0.026$ $0.233\pm 0.009$ $0.262\pm 0.011$ 95%
Electricity $0.165\pm 0.017$ $0.253\pm 0.019$ $0.178\pm 0.002$ $0.270\pm 0.017$ 99%
Traffic $0.416\pm 0.027$ $0.264\pm 0.030$ $0.428\pm 0.008$ $0.282\pm 0.027$ 95%
ETTh1 $0.419\pm 0.023$ $0.432\pm 0.021$ $0.454\pm 0.004$ $0.447\pm 0.007$ 99%
ETTh2 $0.339\pm 0.020$ $0.380\pm 0.019$ $0.383\pm 0.004$ $0.407\pm 0.007$ 95%
ETTm1 $0.369\pm 0.019$ $0.378\pm 0.026$ $0.407\pm 0.004$ $0.410\pm 0.009$ 99%
ETTm2 $0.269\pm 0.021$ $0.320\pm 0.019$ $0.288\pm 0.010$ $0.332\pm 0.003$ 95%
@@
Dataset MAE MAPE RMSE MAE MAPE RMSE Level
@@
Dataset SMAPE MASE OWA SMAPE MASE OWA Level
@@ Appendix H Full Results
Due to the space limitation of the main text, we place the full results of all experiments in the following: long-term forecasting in Table 16, univariate short-term forecasting in Table 17, multivariate short-term forecasting in Table 18, classification in Table 19 and anomaly detection in Table 20.
@@
  • $\ast$ The original paper of Anomaly Transformer (Xu et al., 2022) adopts the temporal association and reconstruction error as a joint anomaly criterion. For fair comparisons, we only use reconstruction error here.